OpenStudy (anonymous):

need help with this algebra problem. Sally can paint a room in 3hrs while it takes steve 6hrs to paint the same room. how long will it take them to paint the room if they work together?

OpenStudy (anonymous):

I don't know if this is right but you can take the terms in terms of the amount of room painted per hour for each person. so Sally can paint 1/3 room per hour, and steve 1/6 room per hour. so to paint one room: y hour x 1/3 room per hour + y hour x 1/6 room per hour = 1 room \[{y \over 3} + {y \over 6} = 1\] and solved for y, the amount of hour they both spent to paint the room.

OpenStudy (anonymous):

There is the first equation y=1-x/3,means how long does it takes from Sally. Thesame to steve y=x/6. Join them together The total task is to finish one room so there is a program of Linear programming just draw a picture,then you will get the answer

OpenStudy (anonymous):

Lol this is the linear algebra section, not algebra.